OpenStudy (anonymous):

Solve for x: 3|x - 3| + 2 = 14

OpenStudy (ankit042):

there are two cases x>3 and x <3 case 1 x> 3 so |x-3| = x-3 3(x-3) +2 = 14 case 2 x <3 so |x-3| = -(x-3) -3(x-3) +2 = 14 Now you can solve!